hello, I can't seem to get the right answer for this problems?
1) A California license plate consists of a number from 1 to 5, then three letters followed by three digits. How many such plates are possible?

2) How many permutations of the letters of the word PROBLEM end in a vowel?

Found 3 solutions by amoresroy, edjones, robertb:
Answer by amoresroy(361)

1)License plates with a number from 1 to 5, then 3 letters, followed by 3 digits
No. of possible combinations = 5 * (26*26*26) * (10*10*10)
= 87,880,000
2) PROBLEM ending in a vowel = 7*6*5*4*3*2 * 2
= 10,080

Answer by edjones(8007)

2)
7!=5040
5040/7=720 end with each letter.
720 end with an O and another 720 end with an E.
1440 Answer.

Answer by robertb(5830)

2) Two ways of filling the rightmost space {o or e}. After that the 6 remaining spaces can be filled in 6! ways, or 720. By the fundamental principle of counting, it can be done in a total of 2*6! = 1440 ways.